Chargent has a capital structure of 40% debt and 60% common equity? The company's tax rate is 25%. Their stock has a beta of 1.4. What is the company's unlevered beta? Calculate the company's levered beta if they changed their capital structure to 20% debt and 80% common equity.
Answer #1

Bu = Bl/(1 + (1 - t)*(D/E))

Bu = 1.4/(1 + 0.75*0.4/0.6)

Bu = 0.93

If they changed their capital structure to 20% debt and 80% common equity

Bu = 1.4/(1 + 0.75*0.2/0.8)

Bu = 1.18